If A= ( -1 2 -3. 4) then e^At

If A = C) 2th 2 4 then eat. = (3e (3et – 2e2t -2et + 2e2t) 3et – 3e2t -2et + 3e2t) de At What is · AeAt dt
Answer #1

(- 4 So detų = 3 2 - Riven H = Here siges Row Sum = 1 is an eigen value as A. We know that dum of eigen So x=1 value=Sum of dAl-Pelp eigen vector Come lo x,-1 i The clor of A(3) is an eigen Vector Coren to Fl. We know that its A is diagonizable ThenLa 2-0 2.0 Vector Corn to 2-2 an eigen So Az = lg 2 iz to G :):):C: 22 :) -2, +aza -32, tuta 22 -32, +42g = 242 - 3 2 -222 =3el zoet So al Potp- 3) ( :) :) 1 2 3 30t -zet ort 30t-reat 20t +2eat 4+382f -2 et +30f So + Ž DR AB+ 31 We know that ex = 1 + 2 + 2 + 3 Al It Alt I (I + Ad + 2 Aga dt = A + d AR + = A (1 + 2A + 2203 ) doab + +